OpenStudy (anonymous):

While simplifying some math work, Peter wrote on his paper that x3 • x3 • x3 • x3 equaled x3+ 3 + 3 +3 . Did Peter simplify his work correctly and completely to a final answer? Would Peter’s work be the same if he were to simplify x3 + x3 + x3 + x3?

OpenStudy (anonymous):

Written Properly: While simplifying some math work, Peter wrote on his paper that \[x ^{3} \times x ^{3} \times x ^{3} \times x ^{3} \] equaled \[x ^{3+3+3+3}\]. Did Peter simplify his work correctly and completely to a final answer? Would Peter’s work be the same if he were to simplify: \[x ^{3} + x^{3} + x ^{3} + x^{3} \] ?

OpenStudy (anonymous):

Let answer be y So the correct answer is y

OpenStudy (anonymous):

You did not answer the question.